# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ VI Agent-Agent Communication
CS Agent Communication Protocol IBM ACP-IBM Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation
Agent Connect Protocol AGNTCY ACP-AGNTCY Allow authenticating callers, threaded state management, flexible execution model
P2P Agent Communication Protocol AgentUnion ACP-AgentUnion Decentralized communication based on the existing domain name system.
Agora Oxford Agora Dynamically switches communication modes based on the communication frequency
Agent Communication Network Fetch.AI ACN Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption.
Agent Network Protocol ANP Team ANP A three-layer architecture and W3C-compliant Decentralized Identifiers.
Layered Orchestration for Knowledgeful Agents CMU LOKA Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance
Hybrid Language Model Operating System Protocol Eclipse LMOS Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management.
Agent to Agent Protocol Google A2A Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms
Others Agent Protocol LangChain Agent Protocol Flexible communication mechanisms based on Run, Thread, and Store.
Agent Interaction & Transaction Protocol NEAR AI AITP Threads-based communication, secure communication across trust boundaries.
@@ VI-A Protocols
We classify the agent-agent communication process into two phases: agent discovery phase and agent communication phase. The first phase is the process in which agents discover the interested agents who satisfy the capability requirement, while the second phase is the task assigning and completing process. According to our analysis, existing protocols show limited differences in the second phase. As a result, we use the first phase as the criterion to classify existing agent-agent communication protocols. Based on it, existing protocols can be divided into four classes: CS-based, Peer-to-peer-based, hybrid, and others (those do not explicitly show their designs in agent discovery).
@@ VI-A1 CS-based Communication
CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions). Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discovery the desired agents. CS-based communication offers stronger agent discovery functionality, such as supporting the search of agents based on capabilities. For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
@@
ACP-IBM. The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [109]. We call it ACP-IBM in this paper to distinguish from the Agent Communication Protocols proposed by other organizations. In ACP-IBM, the client is connected to an agent server. First, the client conducts agent discovery process to discover available agents and get the description of their capabilities. ACP-IBM supports different discover mechanism such as Basic Discovery, Registry-Based Discovery, Offline Discovery, and Open Discovery. Second, after confirming the agent(s), the client start invocation. For single-agent task, the agent server wraps the agent, translating REST calls into internal logic. For multi-agent tasks, the client message is first sent to a Router Agent, which is responsible for decomposing requests, routing tasks, and aggregating responses. ACP-IBM supports synchronous and streaming execution, and allows the preservation of state across multi-turn conversations.
@@
ACP-AGNTCY. The Agent Connect Protocol proposed by AGNTCY [43] is an open standard designed to facilitate seamless communication between agents. The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request. Then, the client further get the agent descriptor by agent ID to know the detailed description of agent functionality. After confirming the target agent, the client can assign tasks to this agent and wait for results. The characteristics of ACP-AGNTCY include flexibility and scalability. First, ACP-AGNTCY deploys a Threads Mechanism, which enables contextual continuity, supporting the creation, copying, and searching of threads, and recording state histories for debugging and backtracking. Second, it supports two operation modes: stateless and stateful. The former is suitable for simple single tasks, while the latter supports multi-round conversations, state continuation and historical data traceability through the thread mechanism to meet the requirements of complex scenarios.
@@ VI-A2 Peer-to-Peer-based Communication
P2P-based communication protocols pursue decentralized agent discovery mechanism. They usually wish use globally universal identifiers (e.g., combined with domain name) to enable agents directly search other agents on the Internet. The advantage of this paradigm is that it supports convenient location and global search (e.g., using crawler) of agents, but they usually do not support the discovery based on agent capability.
ACP-AgentUnion. The Agent Communication Protocol proposed by AgentUnion [5] also aims to provide seamless communication among heterogeneous agents. Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain). As a result, agents can directly access other agents using URI. Agents access IoA through the AP (Access Point), which completes the agent’s identity authentication, address search, communication, and data storage, and provides AID creation, management, and authentication services. In this way, agents can achieve the communication with other agents on the Internet. ACP-AgentUnion does not explicitly support the search of agents based on capabilities.
@@
ACN. Agent Communication Network (ACN) [1, 216] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination. Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels. First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network. Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel, and forward the message after digital signature verification; if not, an error is returned. The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security. Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
ANP (Agent Network Protocol) [249] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents. ANP employs a three-layer architecture. At the Identity and Encrypted Communication Layer, it leverages W3C-compliant Decentralized Identifiers (DIDs) and end-to-end Elliptic Curve Cryptography (ECC) encryption to ensure verifiable cross-platform authentication and confidential agent communication. The Meta-Protocol layer allows agents to dynamically establish and evolve communication protocols through natural language interaction, enabling flexible, adaptive, and efficient inter-agent coordination. At the Application layer, ANP describes agent capabilities using JSON-LD and semantic web standards such as RDF and schema.org, enabling agents to discover and invoke services based on semantic descriptions. It also defines standardized protocol management mechanisms to support efficient and interoperable agent interaction. From a security standpoint, ANP enforces the separation of human authorization from agent-level delegation and adheres to the principle of least privilege. Its minimal-trust, modular design aims to eliminate platform silos and foster a decentralized, composable agent ecosystem.
LOKA. LOKA (Layered Orchestration for Knowledgeful Agents) protocol [217] aims to build a trustworthy and ethical agent ecosystem. Its principle is based on the collaborative operation of multiple key components. First, LOKA introduces the Universal Agent Identity Layer (UAIL), using Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) to assign each agent a unique and verifiable identity, thereby achieving decentralized identity management and autonomous control. Second, LOKA proposes Intent-Centric Communication Protocol, which supports the exchange of semantically rich and ethically annotated messages among agents, promoting semantic coordination and efficient communication. Third, LOKA proposes the Decentralized Ethical Consensus Protocol (DECP). DECP uses multi-party computation (MPC) and distributed ledger technology to enable agents to make context-aware decisions based on a shared ethical baseline, ensuring that their behavior complies with ethical norms. In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
VI-A3 Hybrid Communication
Hybrid communication protocols support both CS-based and P2P-based agent discovery. However, please not that such support is determined by different scenarios. For example, they usually propose a CS-based discovery mechanism specifically for local area network, while the worldwide agent discovery is still P2P-based. In other words, although such protocols support more flexible agent discovery to fit different scenarios, they do not complete eliminate the existing limitations of agent discovery.
LMOS. The LMOS (Language Model Operating System) Protocol proposed by Eclipse [70] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on. LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery. The first method it to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata to the registry. The second method is to use mDNS and DNS-SD protocol to discovery agents/tools in local area networks. The last method is adopting a federal, decentralized protocol (such as P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents. The LMOS protocol has a three-layer architecture. The Application Layer utilizes a JSON-LD-based format to describe the capabilities of agents and tools. The Transport Layer facilitates flexible communication by enabling agents to negotiate protocols like HTTP or MQTT dynamically, accommodating both synchronous and asynchronous data exchange to suit different use cases. The Identity and Security Layer establishes trust through W3C-compliant decentralized identity authentication, combined with encryption and protocols like OAuth2, to secure cross-platform interactions.
A2A. The A2A (Agent to Agent) Protocol proposed by Google [218] aims to enable collaboration between agents. A2A supports three different mechanisms for agent discovery. The first is Well-Known URI, which requires agent servers to store Agent Cards in standardized “well-known” paths uder the domain name (e.g., https://agent-server-domain/.well-known/agent.json). This mechanism enables automatic search of agents on the Internet. However, it does not support the discovery of agents based on capabilities. The second is Curated Registries, i.e., the agents servers register their Agent Cards, which is similar to ACP-IBM. The third is Direct Configuration / Private Discovery, which means that the client can directly requires Agent Cards through hard-coded, local configuration files, environment variables, or private APIs. After finding the desired agents, the client can assign tasks to them and wait for the responses.
VI-A4 Others
This kind of protocols do not explicitly illustrate their unique design for agent discovery. Instead, they only focus on the communication process, e.g., the data format, the management of multiple query, or the historical conversation state.
@@ VI-A4 Others
Agent Protocol. Agent Protocol is proposed by LangChain [142] to enable the communication between LanghGraph (a multi-agent framework) and other types or agents. Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output. Threads act as state containers. They store the cumulative output and checkpoints of multiple rounds of operation. Besides, they support the management of state history (such as querying, copying, and deleting), ensuring that the agent maintains context continuity during multiple rounds of calls. Furthermore, Background Runs support asynchronous task processing and progress can be managed through an independent interface. The element Store provides cross-thread persistent key-value storage for achieving long-term memory. The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks and data storage through HTTP interfaces and configuration parameter. Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
@@ VI-B Security Risk Analysis
Risk Type Attack Detail Attack Characteristics Defense Strategy
CS-based Communication Registration Pollution Malicious registration of fake agents or spam registrations to overload or hijack scheduling Zero-trust registration, frequency/IP monitoring, blacklisting, token-based access (e.g., SAGA)
Description Poisoning Without altering the agent’s identity and by disguising its intended functionality or embedding misleading prompt instructions Capability verification (benchmark+hash), zero-trust registration, control tokens (e.g., SAGA)
Task Flooding Submit a large number of computationally intensive or long-context tasks in a short period, quickly exhausting the resource Load balancing, rate limiting, agent orchestration (e.g., HALO), chaos testing
SEO Poisoning Manipulates agent ranking via keyword stuffing or adversarial optimization Adversarial training, keyword fuzzing, randomized agent selection, historical response quality
P2P-based Communication Non-convergence Lack of central controller leads to repeated or stalled task cycles Lifecycle monitoring, loop detection coordinators, trust graphs (e.g., TMS, G-Memory)
Man-in-the-middle (MITM) Intercepts and tampers messages between agents, exploiting weak encryption or bugs End-to-end encryption (TLS), version updates, redundancy, vulnerability patching
Universal Risks Agent Spoofing Impersonation of trusted agents to inject malicious instructions or steal information Identity authentication (MFA, blockchain DIDs), anomaly detection
Agent Exploitation / Trojan Indirect attack using compromised or malicious helper agents as springboards Agent behavior auditing, sandbox isolation, trace logging (e.g., PeerGuard, CMPL)
Agent Bullying Repeated denial or negative feedback induces cognitive bias or disrupts agent logic Mutual reasoning audits, abnormal feedback pattern detection
Privacy Leakage Unauthorized information spread due to lack of agent-level permission boundaries Fine-grained access control, permission tagging, agent isolation (e.g., PFI, AgentSandbox)
Responsibility Evasion Fault attribution difficult in multi-agent failures or undesired actions Logging, agent-level responsibility tracing (e.g., TRAIL, MSA)
Denial of Service (DoS) Task overloads, prompt loops, or excessive communication drain system resources Agent orchestration, chaos testing, prompt optimization (e.g., HALO, Owotogbe2025)
We make detailed analysis of the security risks in the agent-agent communication process, pointing out possible attacks that have happened and may happen. Since related protocols are getting rapid deployment in various areas, we believe it is urgent to pay more attention to this aspect. We focus more on the structural risks that almost all related protocols will occur instead of tiny design flaws of the existing protocols, which we believe can benefit both the evaluation of the existing deployments and the design of future protocols. In this section, we focus on risks specific for CS-based communication, P2P-based communication, and universal risks for both of them.

### v2

@@ VI Agent-Agent Communication
CS Agent Communication Protocol IBM ACP-IBM[119] Four agent discovery mechanisms, synchronous and streaming execution, multi-turn state preservation
Agent Connect Protocol AGNTCY ACP-AGNTCY[47] Allow authenticating callers, threaded state management, flexible execution model
Agent Communication Protocol AgentUnion ACP-AgentUnion[6] Decentralized APs based on the existing domain name system, each AP holds its agent list
P2P Agent Communication Network Fetch.AI ACN[230] Distributed-Hash-Table-based peer-to-peer discovery, end-to-end encryption.
Agent Network Protocol ANP Team ANP[263] A three-layer architecture and W3C-compliant Decentralized Identifiers.
Layered Orchestration for Knowledgeful Agents CMU LOKA[231] Decentralized identifier, intent-centric communication, privacy-preserving accountability, ethical governance
Hybrid Language Model Operating System Protocol Eclipse LMOS[76] Three agent discovery mechanisms, decentralized digital identifiers, dynamic transport protocol support, group management.
Agent to Agent Protocol Google A2A[89] Three agent discovery mechanisms, cross-platform compatibility, asynchronous priority, security mechanisms
Others Agora Oxford Agora[198] Dynamically switches communication modes based on the communication frequency
Agent Protocol LangChain Agent Protocol[153] Flexible communication mechanisms based on Run, Thread, and Store.
Agent Interaction & Transaction Protocol NEAR AI AITP[9] Threads-based communication, secure communication across trust boundaries.
@@ VI-A Protocols
We classify the agent-agent communication process into two phases: agent discovery phase and agent communication phase. The first phase is the process in which agents discover the interested agents who satisfy the capability requirement, while the second phase is the task assigning and completing process. According to our analysis, existing protocols show limited differences in the second phase. As a result, we use the first phase as the criterion to classify existing agent-agent communication protocols. Based on it, existing protocols can be divided into four classes: CS-based, Peer-to-peer-based, hybrid, and others (those that do not explicitly show their designs in agent discovery).
Figure 7: Agent-agent communication classification: CS-based and P2P-based. Note that the client can also be an agent on the user side.
@@ VI-A1 CS-based Communication
As shown in Figure 7, CS-based communication protocols follow the client-server architecture, which provides centralized servers to manage the information of agents (e.g., their unique IDs and capability descriptions). Under this paradigm, agents interact through well-defined interfaces and rely on centralized servers to discover the desired agents. CS-based communication offers stronger agent discovery functionality, such as supporting the search of agents based on capabilities. For example, the agent servers can run complex search/match algorithms to find proper agent descriptions in their databases.
Figure 8: The communication modes of ACP-IBM. For tasks handled by a single agent (Agent A), the ACP client can directly communicate with it. For tasks requiring multiple agents, a Router Agent acts as a central agent. Note that the client can also be an agent on the user side.
@@
ACP-IBM. The Agent Communication Protocol proposed by IBM is designed for the collaboration of agents [119]. We call it ACP-IBM in this paper to distinguish it from the Agent Communication Protocols proposed by other organizations. In ACP-IBM, the client is connected to an agent server. First, the client conducts an agent discovery process to discover available agents and get the description of their capabilities. ACP-IBM supports different discovery mechanisms such as Basic Discovery, Registry-Based Discovery, Offline Discovery, and Open Discovery. Second, after confirming the agent(s), the client starts the invocation. As shown in Figure 8, for a single-agent task, the agent server wraps the agent, translating REST calls into internal logic. For multi-agent tasks, the client message is first sent to a Router Agent, which is responsible for decomposing requests, routing tasks, and aggregating responses. ACP-IBM supports synchronous and streaming execution, and allows the preservation of state across multi-turn conversations.
@@
ACP-AGNTCY. The Agent Connect Protocol proposed by AGNTCY [47] is an open standard designed to facilitate seamless communication between agents. The client can first search available agents on the agent server, which returns a list of agent IDs matching the criteria provided in the request. Then, the client further gets the agent descriptor by agent ID to know the detailed description of the agent functionality. After confirming the target agent, the client can assign tasks to this agent and wait for results. The characteristics of ACP-AGNTCY include flexibility and scalability. First, ACP-AGNTCY deploys a Threads Mechanism, which enables contextual continuity, supporting the creation, copying, and searching of threads, and recording state histories for debugging and backtracking. Second, it supports two operation modes: stateless and stateful. The former is suitable for simple single tasks, while the latter supports multi-round conversations, state continuation, and historical data traceability through the thread mechanism to meet the requirements of complex scenarios.
ACP-AgentUnion. The Agent Communication Protocol proposed by AgentUnion [6] also aims to provide seamless communication among heterogeneous agents. Each agent has a unique AID (Agent ID), which is a secondary domain name (i.e., agent_name.ap_domain). Agents access IoA through the AP (Access Point), which completes the agent’s identity authentication, address search, communication, and data storage, and provides AID creation, management, and authentication services. As a result, APs can provide the proper agent list based on the query from users. In this way, agents can communicate with other agents on the Internet.
@@ VI-A2 Peer-to-Peer-based Communication
As shown in Figure 7, P2P-based communication protocols pursue a decentralized agent discovery mechanism. They usually wish to use globally universal identifiers (e.g., combined with a domain name) to enable agents to directly search other agents on the Internet. The advantage of this paradigm is that it supports convenient location and global search (e.g., using a crawler) of agents, but they usually do not support the discovery based on agent capability.
ACN. Agent Communication Network (ACN) [230] is a decentralized, peer-to-peer communication infrastructure to facilitate secure and efficient interactions among agents without relying on centralized coordination. Leveraging a Distributed Hash Table (DHT), ACN enables agents to publish and discover public keys, allowing for the establishment of encrypted, point-to-point communication channels. First, agents need to register with one peer node, and the peer node stores the “agent ID - peer node ID” pair in the DHT network. Then, during communication, the source agent sends the message to its associated peer node, and this node recursively searches for the peer node of the target agent through DHT: if the target record exists, the peer nodes of both parties establish a direct communication channel, and forward the message after digital signature verification; if not, an error is returned. The entire communication process uses end-to-end encryption (e.g., TLS) to ensure security. Like the Well-Known URI discovery of A2A, ACN does not support the discovery based on agent capabilities.
ANP (Agent Network Protocol) [263] is an open communication framework designed to enable scalable and secure interoperability among heterogeneous autonomous agents. It supports two types of agent discovery: active and passive. The active discovery uses a uniform URI (.well-known), while the passive discovery submits the agent description to search services. ANP employs a three-layer architecture. At the Identity and Encrypted Communication Layer, it leverages W3C-compliant Decentralized Identifiers (DIDs) and end-to-end Elliptic Curve Cryptography (ECC) encryption to ensure verifiable cross-platform authentication and confidential agent communication. The Meta-Protocol layer allows agents to dynamically establish and evolve communication protocols through natural language interaction, enabling flexible, adaptive, and efficient inter-agent coordination. At the Application layer, ANP describes agent capabilities using JSON-LD and semantic web standards such as RDF and schema.org, enabling agents to discover and invoke services based on semantic descriptions. It also defines standardized protocol management mechanisms to support efficient and interoperable agent interaction. From a security standpoint, ANP enforces the separation of human authorization from agent-level delegation and adheres to the principle of least privilege. Its minimal-trust, modular design aims to eliminate platform silos and foster a decentralized, composable agent ecosystem.
LOKA. LOKA (Layered Orchestration for Knowledgeful Agents) protocol [231] aims to build a trustworthy and ethical agent ecosystem. Its principle is based on the collaborative operation of multiple key components. First, LOKA introduces the Universal Agent Identity Layer (UAIL), using Decentralized Identifiers (DIDs) and Verifiable Credentials (VCs) to assign each agent a unique and verifiable identity, thereby achieving decentralized identity management and autonomous control. Second, LOKA proposes an Intent-Centric Communication Protocol, which supports the exchange of semantically rich and ethically annotated messages among agents, promoting semantic coordination and efficient communication. Third, LOKA proposes the Decentralized Ethical Consensus Protocol (DECP). DECP uses multi-party computation (MPC) and distributed ledger technology to enable agents to make context-aware decisions based on a shared ethical baseline, ensuring that their behavior complies with ethical norms. In addition, the authors also point out that it combines cutting-edge technologies such as distributed identity, verifiable credentials, and post-quantum cryptography to provide comprehensive support for the agent ecosystem in terms of identity management, communication and coordination, ethical decision-making, and security.
VI-A3 Hybrid Communication
Hybrid communication protocols support both CS-based and P2P-based agent discovery. However, please note that such support is determined by different scenarios. For example, they usually propose a CS-based discovery mechanism specifically for local area networks, while the worldwide agent discovery is still P2P-based. In other words, although such protocols support more flexible agent discovery to fit different scenarios, they do not completely eliminate the existing limitations of agent discovery.
LMOS. The LMOS (Language Model Operating System) Protocol proposed by Eclipse [76] aims to enable agents and tools from diverse organizations to be easily discovered and connected, regardless of the technologies they are built on. LMOS supports three different agent discovery methods to enable both centralized and decentralized discovery. The first method is to adopt the mechanism of W3C Web of Things (WoT) to enable agents to dynamically register metadata in the registry. The second method is to use mDNS and DNS-SD protocol to discover agents/tools in local area networks. The last method is adopting a federal, decentralized protocol (such as P2P protocol) to disseminate agents and tool descriptions, without relying on a centralized registry center, which is applicable for global collaboration of agents. The LMOS protocol has a three-layer architecture. The Application Layer utilizes a JSON-LD-based format to describe the capabilities of agents and tools. The Transport Layer facilitates flexible communication by enabling agents to negotiate protocols like HTTP or MQTT dynamically, accommodating both synchronous and asynchronous data exchange to suit different use cases. The Identity and Security Layer establishes trust through W3C-compliant decentralized identity authentication, combined with encryption and protocols like OAuth2, to secure cross-platform interactions.
A2A. The A2A (Agent to Agent) Protocol proposed by Google [89] aims to enable collaboration between agents. A2A supports three different mechanisms for agent discovery. The first is Well-Known URI, which requires agent servers to store Agent Cards in standardized “well-known” paths under the domain name (e.g., https://agent-server-domain/.well-known/agent.json). This mechanism enables automatic search of agents on the Internet. However, it does not support the discovery of agents based on capabilities. The second is Curated Registries, i.e., agent servers register their Agent Cards, which is similar to ACP-IBM. The above two methods can be referred to as Figure 7. The third is Direct Configuration / Private Discovery, which means that the client can directly require Agent Cards through hard-coded, local configuration files, environment variables, or private APIs. After finding the desired agents, the client can assign tasks to them and wait for the responses.
VI-A4 Others
This kind of protocols do not explicitly illustrate their unique design for agent discovery. Instead, they only focus on the communication process, e.g., the data format, the management of multiple queries, or the historical conversation state.
@@ VI-A4 Others
Agent Protocol. Agent Protocol is proposed by LangChain [153] to enable the communication between LanghGraph (a multi-agent framework) and other types of agents. Its mechanism is based on Thread and Run: Run is a single call of the agent, which supports streaming output of real-time results or waiting for the final output. Threads act as state containers. They store the cumulative output and checkpoints of multiple rounds of operation. Besides, they support the management of state history (such as querying, copying, and deleting), ensuring that the agent maintains context continuity during multiple rounds of calls. Furthermore, Background Runs support asynchronous task processing, and progress can be managed through an independent interface. The element Store provides cross-thread persistent key-value storage for achieving long-term memory. The overall mechanism realizes flexible control over proxy calls, status management, asynchronous tasks and data storage through HTTP interfaces and configuration parameters. Agent Protocol does not explicitly illustrate the unique agent discovery mechanism it supports.
@@ VI-B Security Risk Analysis
Risk Type Representative Threats Attack Characteristics
In CS-based Communication Registration Pollution Malicious registration of fake/nonexistent agents
Description Poisoning Embed malicious prompt instructions in the description
Task Flooding Submit a large number of computationally intensive or long-context tasks
SEO Poisoning Manipulates agent ranking via keyword stuffing or adversarial optimization
In P2P-based Communication Non-convergence The execution of a task falls into a loop, unable to reach the finishment
Man-in-the-middle (MITM) Intercept, tamper with, or replay messages between agents
Universal Risks Agent Spoofing Impersonation of trusted agents to inject malicious instructions
Agent Exploitation / Trojan Indirect attack using compromised or malicious helper agents as springboards
Agent Bullying Repeated denial, negative, or aggressive feedback to benign agents
Privacy Leakage Unauthorized information spread due to the lack of permission boundaries
Responsibility Evasion Fault attribution is difficult in multi-agent failures or undesired actions
Denial of Service (DoS) Task overloads, prompt loops, or excessive communication drain system resources
We make a detailed analysis of the security risks in the agent-agent communication process, pointing out possible attacks that have happened and may happen. Since related protocols are getting rapid deployment in various areas, we believe it is urgent to pay more attention to this aspect. We focus more on the structural risks that almost all related protocols will encounter instead of the tiny design flaws of the existing protocols, which we believe can benefit both the evaluation of the existing deployments and the design of future protocols. In this section, we focus on risks specific to CS-based communication, P2P-based communication, and universal risks for both of them.